OpenStudy (anonymous):

Create your own third degree polynomial that when divided by x + 2 has a remainder of –4

OpenStudy (cwrw238):

(-2)^3 + 4(-2)^2 + (-2) -2 = 4 so its x^3 + 4x^2 + x - 2

OpenStudy (cwrw238):

i used the Remainder Theorem: if f(x) is divided by (x + a) then remainder is f(-a)

OpenStudy (cwrw238):

oh - sorry mape - misread the question - the remainder must be - 4 i'll redo it

OpenStudy (cwrw238):

we f(-2) to be = -4 (-2)^3 + 4(-2)^2 + (-2) = -8 + 16 -2 = 6 so to get -4 we need to subtract - 10 so the polynomial will be x^3 + 4(x)^2 + x - 10

OpenStudy (cwrw238):

ok?

OpenStudy (anonymous):

i wonder, can it be any polynomial: (x+2)(ax^2 + bx + c) - 4

OpenStudy (cwrw238):

yes it can

OpenStudy (anonymous):

yes, using remainder theorem f(-2) = -4 so if f(x) = (x+2)(ax^2 + bx + c) - 4 for all a,b,c f(-2) = -4

OpenStudy (cwrw238):

good thinking

OpenStudy (anonymous):

your answer is just as valid